Problem 6 Calculate the test statistic. (A z-score and a t-score are examples of test statistics.) In order to present your final results with a numerical accuracy of 0.01, it will be necessary to perform your intermediate calculations with an accuracy of 0.001. Part 1 (696) The cost of a daily newspaper varies from city to city. A study was done to test the claim that the average cost of a daily newspaper is $0.40 Fifteen cities were surveyed; the average cost was $0.50 with a standard deviation of $0.05 Part 2 (6%) Your statistics instructor claims that 60 percent of the students who take an Elementary Statistics class go through life feeling more enriched. For some mysterious reason, most people dont believe this. You decide to check this out on your own、 You randomly survey 80 former Elementary Statistics students and find that 30 feel more enriched as a result of the class
Answer #1

part 1) test statistic t=(xbar-mean)*sqrt(n)/std deviation=(0.5-0.4)*sqrt(15)/0.05=7.75

part 2)here sample proportion phat=30/80=0.375

std error =sqrt(p(1-p)/n)=0.0548

hence z =(phat-p)/std error =(0.375-0.6)/0.0548=-4.11